OpenStudy (anonymous):

by foil method: (2m-5n) (3n-7n)

OpenStudy (mimi_x3):

6mn - 14mn - 15n^2 + 35n^2 6mn - 14mn +20n^2

OpenStudy (anonymous):

i need other answer pls

OpenStudy (mimi_x3):

other answer ?

OpenStudy (anonymous):

yes

OpenStudy (mimi_x3):

but there's only one answer to this question

OpenStudy (anonymous):

by horizontal method: (2m-5n) (3m-7n)

OpenStudy (mimi_x3):

oh horizontal method!? I dont know what a horizontal method sorry

OpenStudy (anonymous):

http://www.purplemath.com/modules/polymult.htm

OpenStudy (mimi_x3):

is horizontal method like foil method ?

OpenStudy (anonymous):

(sm-5n)(3m-7n) =(2m-5n)3m - (2m-5n)7n and so on ...

OpenStudy (anonymous):

is this a answer

OpenStudy (mimi_x3):

oh but you would get the same answer anyway

OpenStudy (anonymous):

you would, if you didn't mess up :) (2m-5n)(2m-7n) =(2m-5n)3m - (2m-5n)7n =6m^2-15mn - 14mn+35n^2 =6m^2 -29mn +35n^2

OpenStudy (anonymous):

which I do a lot ...

OpenStudy (mimi_x3):

looks complicated , the foil method is better xD

OpenStudy (anonymous):

I like box method best !!

OpenStudy (mimi_x3):

what's box method ?

OpenStudy (anonymous):